(Expected rate of return and​ risk) ​ Syntex, Inc. is considering an investment in one of two common stocks. Given the information that​ follows, which investment is​ better, based on the risk​ (as measured by the standard​ deviation) and​ return? Common Stock A              Common Stock B              Probability Return Probability Return 0.25 13​% 0.10 negative 7​% 0.50 17​% 0.40 8​% 0.25 20​% 0.40 13​% 0.10 20​%
Answer #1

A:
Expected returns=0.25*13%+0.50*17%+0.25*20%=16.75000%

Standard deviation=sqrt(0.25*(13%-16.75%)^2+0.50*(17%-16.75%)^2+0.25*(20%-16.75%)^2)=2.48747%

B:
Expected returns=0.10*(-7%)+0.40*8%+0.40*13%+0.10*20%=9.70000%

Standard deviation=sqrt(0.10*(-7%-9.7%)^2+0.40*(8%-9.7%)^2+0.40*(13%-9.7%)^2+0.10*(20%-9.7%)^2)=6.63400%

Stock A is better